OpenStudy (anonymous):

Solve for the unknown variable: please help! 1) 2(x+1) = (3x-2)+1 2) y+3/4-y+1/2=-4 thanks guys..

OpenStudy (anonymous):

2(x+1) = (3x-2) + 1 2x + 2 = 3x - 1 (Distribution) 2x = 3x -3 (Subtract 2 from both sides) -x = -3 (Multiply both sides by -1 x = 3

hero (hero):

yup

OpenStudy (anonymous):

thanks woufon. can u help with the 2nd one as well